OpenStudy (anonymous):

∫(5x^3+ 4x-8)dx - evaluate

OpenStudy (angela210793):

=5x^4/4+4x^2/2-8x+c

OpenStudy (anonymous):

how did you get ^ 4

OpenStudy (angela210793):

integral of x^n=x^(n+1)/(n+1)

OpenStudy (anonymous):

it's the definition of the antiderivative

OpenStudy (anonymous):

what about the c and 8x?

OpenStudy (anonymous):

c is the constant you add after every indefinite integral the integral of a constant is constant * x

OpenStudy (anonymous):

its necessary to put a constant?

OpenStudy (angela210793):

well every constant has an integral =c-te*x and u add +c cause it is a indefinite integral

OpenStudy (angela210793):

yes whn u have an indefinite integral

OpenStudy (anonymous):

how can you distinguish wether its indefinite and definite

OpenStudy (anonymous):

It's necessary if it's an indefinite integral. When you have a definite integral (an integral with a lower and an upper bound included) you do not have to consider the c

OpenStudy (angela210793):

a definite integral is \[\int\limits_{a}^{b}\]

OpenStudy (anonymous):

\[\int\limits_{0}^{2}xdx\] is a definite integral with a lower bound of 0 and an upper bound of 2

OpenStudy (anonymous):

any reason or anywhere integrals can come of use in work or in professional fields?

OpenStudy (angela210793):

well definite integrals are used to find areas

OpenStudy (angela210793):

http://en.wikipedia.org/wiki/Integral take a look

OpenStudy (anonymous):

thankx

OpenStudy (angela210793):

yw :)